# Origin of Pulsed Radio Emission from Magnetars

### v1

@@ Appendix B Prediction of saturated electric field energy: a quasi-linear argument
The strength of the synchrotron cooling can be parametrized by comparing the timescale of synchrotron losses $t_{\rm syn}$ and local light crossing time, $t_{\rm syn}/(r/c)=3/(4\gamma\mathscr{L}_{B}\sin^{2}\alpha)$, where
@@
is the magnetic compactness parameter, $\alpha$ is the particle’s pitch angle, $U_{B}=B^{2}/8\pi$ is the energy density of magnetic field, $\sigma_{T}$ is the Thomson cross section. In the inner magnetosphere, $r\lesssim 50R_{*}$, synchrotron losses occur on a timescale much shorter than the timescale of the radiative drag, $t_{\rm syn}\lesssim 10^{-7}(r/c)\sim 10^{-6}t_{\rm cool}$, and can therefore be considered instantaneous. The synchrotron cooling timescale becomes comparable to the light crossing time at $r\sim 10^{3}R_{*}$ due to the rapid, $\propto r^{-3}$, decay of the magnetic field strength.
@@
Lyutikov (2020) considered the same dispersion relation and identified unstable branches. Their interpretation of the instability near $k=0$ as a firehose instability, however, is not correct. In fact, by rewriting the instability threshold as $k<8\pi nev_{d}/(Bc)=4\pi j_{0}/(Bc)$, it is clear that this is a kink-like instability relevant only on large spatial scales, $\lambda>Bc/(2j_{0})\approx(2\pi/\psi)R_{\mathrm{max}}$, and is therefore irrelevant for quiescent stable magnetospheres.
@@ Appendix F Optical Depth of cyclotron absorption
The cyclotron absorption can occur when the frequency of the electromagnetic waves seen by the particles matches with the cyclotron frequency, $\gamma\tilde{\omega}=\gamma(1-\beta\cos\phi)\omega=\omega_{B}$, where $\phi$ is the angle between the wave propagation direction and the particle velocity. Taking $\nu\sim 10\mathrm{GHz}$ and assuming that the effect of the Doppler shift is of order unity, it can be estimated that the cyclotron resonance occurs at $r_{c}\sim$ several $\times 10^{3}R_{*}$. Since the resonance occurs far from the magnetar, we can make the approximation that the emission propagates radially. The cross-section for cyclotron absorption is $\sigma_{\rm cyc}=2\pi^{2}e^{2}/(m_{e}c)\delta(\gamma\tilde{\omega}-\omega_{B})$, and the optical depth of the radiation emitted in the direction along the field line, at $(r_{0},\theta_{0})$, is
@@ Appendix G Polarization of Normal modes
Polarization of normal modes at large altitudes where radio emission decouples from the plasma can be calculated by solving the dispersion relation for the dielectric tensor corresponding to Eq. (C2). The difference between the refractive indexes of two electromagnetic normal modes can be expressed as (e.g., Cheng & Ruderman, 1979; Beskin & Philippov, 2012):

### v2

@@ Appendix B Prediction of saturated electric field energy: a quasi-linear argument
The strength of the synchrotron cooling can be parametrized by comparing the timescale of synchrotron losses $t_{\rm syn}$ and local light crossing time, $t_{\rm syn}/(r/c)=3/(4\gamma\mathscr{L}_{B}\sin^{2}\chi)$, where
@@
is the magnetic compactness parameter, $\chi$ is the particle’s pitch angle, $U_{B}=B^{2}/8\pi$ is the energy density of magnetic field, $\sigma_{T}$ is the Thomson cross section. In the inner magnetosphere, $r\lesssim 50R_{*}$, synchrotron losses occur on a timescale much shorter than the timescale of the radiative drag, $t_{\rm syn}\lesssim 10^{-7}(r/c)\sim 10^{-6}t_{\rm cool}$, and can therefore be considered instantaneous. The synchrotron cooling timescale becomes comparable to the light crossing time at $r\sim 10^{3}R_{*}$ due to the rapid, $\propto r^{-3}$, decay of the magnetic field strength.
@@
lyutikov2020 considered the same dispersion relation and identified unstable branches. Their interpretation of the instability near $k=0$ as a firehose instability, however, is not correct. In fact, by rewriting the instability threshold as $k<8\pi nev_{d}/(Bc)=4\pi j_{0}/(Bc)$, it is clear that this is a kink-like instability relevant only on large spatial scales, $\lambda>Bc/(2j_{0})\approx(2\pi/\psi)R_{\mathrm{max}}$, and is therefore irrelevant for quiescent stable magnetospheres.
@@ Appendix F Optical Depth of cyclotron absorption
The cyclotron absorption can occur when the frequency of the electromagnetic waves seen by the particles matches with the cyclotron frequency, $\gamma\tilde{\omega}=\gamma(1-\beta\cos\phi)\omega=\omega_{B}$. Taking $\nu\sim 10\mathrm{GHz}$ and assuming that the effect of the Doppler shift is of order unity, it can be estimated that the cyclotron resonance occurs at $r_{c}\sim$ several $\times 10^{3}R_{*}$. Since the resonance occurs far from the magnetar, we can make the approximation that the emission propagates radially. The cross-section for cyclotron absorption is $\sigma_{\rm cyc}=2\pi^{2}e^{2}/(m_{e}c)\delta(\gamma\tilde{\omega}-\omega_{B})$, and the optical depth of the radiation emitted in the direction along the field line, at $(r_{0},\theta_{0})$, is
@@ Appendix G Polarization of Normal modes
Polarization of normal modes at large altitudes where radio emission decouples from the plasma can be calculated by solving the dispersion relation for the dielectric tensor corresponding to Eq. (C2). The difference between the refractive indexes of two electromagnetic normal modes can be expressed as (e.g., ChengRuderman1979; BeskinPh2012):